Multiply both sides of the equation: 9x divided by 4 plus 3 equals 1 eighth by 8. What does the function "by" mean and how do you solve this?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Multiply both sides of the equation: 9x divided by 4 plus 3 equals 1 eighth by 8. What does the function "by" mean and how do you solve this?
----------
What you posted is not clear.
---------
"Divided by" is, eg, a "divided by" b = a/b
3 divided by 4 = 3/4 = 0.75 decimal.
---------
1 divided by 2 = 1/2, one half.
etc.

Answer by ikleyn(52776)

.
Read this Wikipedia article
https://en.wikipedia.org/wiki/Division_(mathematics)

https://en.wikipedia.org/wiki/Division_(mathematics)

especially its chapter called "Notation".